# MERaLiON-AudioLLM: Technical Report

### v1

@@ 2.1.1 Audio Encoder
We are also exploring the integration of a localised speech encoder, which has been pre-trained from scratch using a self-supervised learning (SSL) framework, and will introduce it in future work.
@@ 5.1 Setup
  • ASR: LibriSpeech (Panayotov et al., 2015), Common-Voice-15 (Ardila et al., 2020), Earnings21 (Rio et al., 2021), Earnings22 (Rio et al., 2022), NSC (Koh et al., 2019)
@@
  • SQA: SLUE (Shon et al., 2023), Spoken-SQuAD (Lee et al., 2018), CN-College-Listen-Test (Hu et al., 2024), Singapore-Public-Speech-SQA (Wang et al., 2024), NSC (Koh et al., 2019)
  • SDS: NSC (Koh et al., 2019)
@@ 5.2 Results
MNSC-ASR-Part 1 0.05 0.07 - 0.09 0.07
MNSC-ASR-Part 2 0.05 0.19 - 0.42 0.33
MNSC-ASR-Part 3 0.28 0.35 - 0.66 0.30
MNSC-ASR-Part 4 0.40 0.56 - 0.76 0.48
MNSC-ASR-Part 5 0.21 0.28 - 0.35 0.23
MNSC-ASR-Part 6 0.15 0.22 - 0.25 0.18
@@
CoVoST 2 En $\rightarrow$ Id 32.62 16.33 13.84 14.14 27.62
CoVoST 2 En $\rightarrow$ Zh 37.98 25.77 31.96 33.89 35.27
CoVoST 2 En $\rightarrow$ Ta 8.50 0.03 0.00 0.00 8.46
CoVoST 2 Id $\rightarrow$ En 37.07 6.33 5.93 26.89 46.80
CoVoST 2 Zh $\rightarrow$ En 15.01 16.47 2.37 5.30 15.21
CoVoST 2 Ta $\rightarrow$ En 3.97 0.04 0.17 0.36 2.83
@@
SLUE-SQA-5 82.94 80.05 83.92 83.48 88.58
Spoken-SQuAD 70.33 64.86 77.65 66.40 88.62
CN-College-Listen-Test 85.03 74.51 65.43 50.90 91.85
Singapore-Public-Speech-SQA 60.32 58.31 58.55 59.24 73.11
MNSC-SQA-Part 3 51.4 42.0 - 40.60 53.20
MNSC-SQA-Part 4 49.0 39.6 - 36.60 60.20
MNSC-SQA-Part 5 58.2 51.6 - 44.60 67.20
MNSC-SQA-Part 6 65.2 53.6 - 46.80 71.60
@@
MNSC-SDS-Part 3 46.80 33.80 - 9.0 45.40
MNSC-SDS-Part 4 45.80 24.80 - 7.0 44.00
MNSC-SDS-Part 5 55.2 40.4 - 17.2 58.00
MNSC-SDS-Part 6 61.8 46.2 - 24.2 65.40
@@
OpenHermes-Audio 71.4 44.8 22.40 15.80 72.20
Alpaca-GPT4-Audio 73.4 52.6 21.60 17.20 73.80
@@
VoxCeleb-Gender-Test 99.53 99.12 69.68 88.81 35.25
VoxCeleb-Accent-Test 46.35 29.18 - 34.22 24.64
MELD-Sentiment-Test 42.26 53.49 50.08 42.07 56.67
MELD-Emotion-Test 30.15 40.54 41.07 30.73 47.39

### v3

# MERaLiON-AudioLLM: Bridging Audio and Language with Large Language Models

@@ 2.1.1 Audio Encoder
We are also exploring the integration of a localised speech encoder, which has been pre-trained from scratch using a self-supervised learning (SSL) framework, and will introduce it in future work MERaLiON Team (2024).
@@ 5.1 Setup
  • ASR: LibriSpeech (Panayotov et al., 2015), Common-Voice-15 (Ardila et al., 2020), Earnings21 (Rio et al., 2021), Earnings22 (Rio et al., 2022), NSC (Koh et al., 2019; Wang et al., 2025)
@@
  • SQA: SLUE (Shon et al., 2023), Spoken-SQuAD (Lee et al., 2018), CN-College-Listen-Test (Hu et al., 2024), Singapore-Public-Speech-SQA (Wang et al., 2024), NSC (Koh et al., 2019; Wang et al., 2025)
  • SDS: NSC (Koh et al., 2019; Wang et al., 2025)
@@ 5.2 Results
MNSC-ASR-Part 1 0.05 0.07 0.10 0.09 0.07
MNSC-ASR-Part 2 0.05 0.19 0.45 0.42 0.33
MNSC-ASR-Part 3 0.28 0.35 0.75 0.66 0.30
MNSC-ASR-Part 4 0.40 0.56 1.14 0.76 0.48
MNSC-ASR-Part 5 0.21 0.28 0.40 0.35 0.23
MNSC-ASR-Part 6 0.15 0.22 0.43 0.25 0.18
@@
CoVoST 2 En $\rightarrow$ Id 32.6 16.3 13.8 14.1 27.6
CoVoST 2 En $\rightarrow$ Zh 38.00 25.8 32.0 33.9 35.3
CoVoST 2 En $\rightarrow$ Ta 8.5 0.0 0.0 0.0 8.5
CoVoST 2 Id $\rightarrow$ En 37.1 6.3 5.9 26.9 46.8
CoVoST 2 Zh $\rightarrow$ En 15.0 16.5 2.4 5.3 15.2
CoVoST 2 Ta $\rightarrow$ En 4.0 0.0 0.2 0.4 2.8
@@
SLUE-SQA-5 82.9 80.1 83.9 83.5 88.6
Spoken-SQuAD 70.3 64.9 77.7 66.4 88.6
CN-College-Listen-Test 85.0 74.5 65.4 50.9 91.9
Singapore-Public-Speech-SQA 60.3 58.3 58.6 59.2 73.1
MNSC-SQA-Part 3 51.4 42.0 45.2 40.6 53.2
MNSC-SQA-Part 4 49.0 39.6 46.6 36.6 60.2
MNSC-SQA-Part 5 58.2 51.6 50.8 44.6 67.2
MNSC-SQA-Part 6 65.2 53.6 62.2 46.8 71.6
@@
MNSC-SDS-Part 3 46.8 33.8 31.6 9.0 45.4
MNSC-SDS-Part 4 45.8 24.8 31.6 7.0 44.0
MNSC-SDS-Part 5 55.2 40.4 45.2 17.2 58.0
MNSC-SDS-Part 6 61.8 46.2 49.4 24.2 65.4
@@
OpenHermes-Audio 71.4 44.8 22.4 15.8 72.2
Alpaca-GPT4-Audio 73.4 52.6 21.6 17.2 73.8
@@
VoxCeleb-Gender-Test 99.5 99.1 69.7 88.8 35.3
VoxCeleb-Accent-Test 46.4 29.2 35.8 34.2 24.6
MELD-Sentiment-Test 42.3 53.5 50.1 42.1 56.7
MELD-Emotion-Test 30.2 40.5 41.1 30.7 47.4
@@ Acknowledgement
We extend our sincere gratitude to Jeremy H. M. Wong, Kye Min Tan, Geyu Lin, Ziyi Xu, Hardik B. Sailor, Qiongqiong Wang, Tianchi Liu, Muhammad Huzaifah, Xin Huang, Tarun K. Vangani, Nattadaporn Lertcheva, Xi Wang, Kui Wu, Siti Maryam Binte Ahmad Subaidi, Madalene Hee for their invaluable contributions to data management, human annotations, logistics, insightful discussions, and future work explorations.